OpenStudy (anonymous):

Find the particular solution to y ' = 2sin(x) given the general solution is y = C − 2cos(x) and the initial condition y of pi over 2 equals 1.

OpenStudy (irishboy123):

what do you think?? it all seems to be included in the info pack.....

OpenStudy (anonymous):

What info pack?

OpenStudy (irishboy123):

you have the DE and the solution. what is there left to do apart from plug and play!!

OpenStudy (anonymous):

Can you show me how it's done? I have about 3 more of the same question and I'm not really sure how to approach it. What do I plug and why?

OpenStudy (freckles):

you have the condition y(pi/2)=1 this says when x=pi/2 we have y=1 plug into your solution to find C

OpenStudy (freckles):

do you see your solution is y=C-2cos(x) right? Find the constant C using the point (x=pi/2,y=1) replace x with pi/2 replace y with 1 solve for C you can do this! :)

OpenStudy (anonymous):

C=1 Final answer is C?

OpenStudy (freckles):

\[y=C-2 \cos(x) \\ 1=C-2 \cdot \cos(\frac{\pi}{2}) \\ 1=C-2(0) \\1=C \\ \text{ yes you should have } \\ y=1-2\cos(x)\]